# reference.bib

### v2

@@ 5.3.3 Decomposition Benefits
Figure 3: Monetary ratio of strategyQA and MuSiQue dataset before (decomp) and after (w/o decomp) decompositon under various privacy budget $\mu$. Decompostion in ConfusionPrompt significantly reduces the monetary cost, indicating its efficiency.
We evaluate generation on longer attributes, memory cost of local models, the performance of the discriminator $D$ that verifies the genuineness, and pseudo prompt generation cost (see Appendix F.5, F.7, F.8, and F.9).
6 Conclusion
This paper proposes a private inference framework over online LLMs, termed ConfusionPrompt. We deploy three local models on the user side: (i) decomposer that maps an original prompt to a sequence of sub-prompts, (ii) generator that produces pseudo prompts by replacing private attributes in the genuine sub-prompts, (iii) recomposer that maps the decomposed prompt-response pairs from the cloud service to the final response. Such designs endows our framework with advantages over previous protocols that: (i) it can be seamlessly integrated with existing black-box LLMs while the service providers have no need to modify their already-built framework, and (ii) it achieves better privacy-utility trade-off than existing privatization methods. We develop a $(\lambda,\mu,\rho)$-privacy model to formulate the requirement for a privacy-preserving group of prompts, and accordingly provide a complexity analysis to demonstrate the benefits of the decomposition module. Experiments show that our ConfusionPrompt significantly outperforms the inference with local open-sourced LLMs in terms of utility and computation cost.
[…]
Appendix A Decomposition Example
Figure 4 presents an example demonstrating the decomposition benefits.
Figure 4: Example of decomposition savings in query complexity. Decomposition module reduces the query complexity from 9 to 6 under privacy budget $\mu=3$.
Appendix B Proof of Theorem 4.1 and 4.2
We begin with the proof for Theorem 4.1 as followed:
Proof.
$\begin{gathered}\mathrm{Sig}(U(p_{0}),\mathbb{P})=\max_{u_{k}}\max_{\mathcal{V}}\frac{\sum_{i=1}^{n}\bar{H}(\mathcal{V},p_{0},p_{i})}{\sum_{i=1}^{n}\bar{H}(\mathcal{V}\backslash u_{k},p_{0},p_{i})}\\ {\rm s.t.}u_{k}\in\mathcal{V}\subseteq U(p_{0}))\end{gathered}$ (10)
$\bar{H}(\mathcal{V},p_{0},p_{i})=\left\{\begin{array}[]{ll}1&\mathrm{Corr}\left(u,p_{0},p_{i}\right)=u\ \forall u\in\mathcal{V}\\ &\mathrm{or}\ \mathcal{V}=\emptyset\\ 0&\rm{otherwise}\end{array}\right.$ (11)
For $\mathrm{Sig}(U(p_{0}),\mathbb{P})=\mu$, it’s obvious that there should be at least $1/\mu$ distinct values, including genuine and fake values, assigned to each attributes. Otherwise suppose $u_{k}$ have less than $1/\mu$ distinct values, we would have:
$\frac{\sum_{j=1}^{|\mathbb{P}|}H(u_{k},p_{0},p_{j})}{|\mathbb{P}|}>\mu$ (12)
which violates the expression 10.
(b) The missing combination contain a subset of true values $\mathcal{V}\subseteq U(p_{0})$. Then conditioning on the subset $\mathcal{V}$, the proportion of any other single attribute would be over $\mu$, as their original proportion is $\mu$. Therefore, for any $u_{k}\notin\mathcal{V}$, it holds that:
$\frac{\sum_{i=1}^{n}\bar{H}(\mathcal{V}\cup u_{k},p_{0},p_{i})}{\sum_{i=1}^{n}\bar{H}(\mathcal{V},p_{0},p_{i})}>\mu$ (13)
which violates the expression 10.
Following that, we can proceed to the proof for Theorem 4.2:
Proof.
Let $\mu=\mathrm{Sig}(U(p_{0}),\mathbb{P})$ denotes the significance. According to Theorem 4.1, for each sub-query $p_{0h}\in\boldsymbol{p}_{0}$, we should have at least $(1/\mu)^{\|U(p_{0h})\|}$ prompts to satisfy:
$\max_{u_{k}}\\ \max_{\mathcal{V}}\frac{\sum_{i=1}^{n}\bar{H}(\mathcal{V},p_{0h},p_{ih})}{\sum_{i=1}^{n}\bar{H}(\mathcal{V}\backslash u_{k},p_{0h},p_{ih})}\leq\mu\quad\quad{\rm s.t.}u_{k}\in\mathcal{V}\subseteq U(p_{0h})$ (14)
Appendix C Bound on Inference Attack
In this section, we consider the attacker identify the target attribute by prior knowledge of other attributes and the combination pattern of attribute values. The following theorem states that an curious server can do no better than random guessing if the combination of values are uniformly distributed:
Lemma C.1.
Suppose the attacker is guessing the attributes from attribute combination pattern in $\mathbb{P}$. The attacker can do no more better than random guessing if the following condition is satisfied:
  • Each combination of attribute set has the same occurrence times in the prompt groups $\mathbb{P}$, i.e., the combinations are uniformly distributed.
From the proof of Theorem 4.1 and 4.2, we can construct a uniformly distributed prompt group under the optimal number of prompts. Therefore, we can turn to the success rate under an random guessing attack, which is bounded by Lemma C.2.
Lemma C.2.
Given a group of prompts =$\mathbb{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ with $(\lambda,\mu,\rho)$-privacy, suppose that the attacker have prior knowledge about the attribute subset $\mathcal{V}\subseteq U(\boldsymbol{p}_{0})$. By random guessing, the probability of correctly identifying any target attribute $u\in U(\boldsymbol{p}_{0})\land u\notin\mathcal{V}$ is upper bounded by $\mu$.
Proof.
According to Definition 3.8, it holds that:
Appendix D Ranking Rules in Preference Sampling
D.1 Decomposer
We rank the decompositions with the following steps:
  • Identify whether or not each decomposition satisfies the MECE principle.
  • Compute the complexity of each decompositon under a given $\mu$.
  • Rank the responses using the following rules: (i) Decompositions that adhere to the MECEprinciple should be ranked higher than those that do not follow the principle. (ii) If two decompositions both follow or do not follow the MECE principle, the one with lower complexity is preferred.
D.2 Generator
To rank the pseudo prompts, we take the steps below:
  • Compute $s$, the similarity between the pseudo and original prompt.
  • Compute $f$, the genuineness score for the pseudo prompt.
  • Rank the responses with the overall score, with a higher score indicating a higher preference.
Appendix E Algorithm
In this section we present the algorithm of ConfusionPrompt in 1.
Algorithm 1 ConfusionPrompt
2:  Output: Final response $r$.
4:  for $h\in[|\boldsymbol{p}_{0}|]$ do
5:     User keeps sampling from the generator until there are $n$ distinct pseudo prompts with $n$ given by eq. 7, where each one $1\leq i\leq n$ has prompt similarity $\mathrm{Sim}(p_{0h},p_{ih})\leq\lambda$ and genuineness score $\mathrm{Genu}(p_{ih})\geq\rho$.
6:  end for
7:  User sends the group of genuine and pseudo prompts =$\mathbb{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ to the server.
8:  Server returns a collection of responses $(r_{0},...,r_{l})$.
9:  User retrieves the response corresponding to the genuine prompt $r_{0}$.
Appendix F Experiment
F.1 Construction of Private Attributes
We extract the private attributes of StrategyQA and MuSiQue through the steps below:
  • Step 1: sample construction. We adopt a semi-automated method to extract private attributes from each sample. Firstly, we extract the entities using a combination of two NER methods: Spacy [srinivasa2018natural] and Flair [akbik2019flair]. Then we manually correct and supplement the private attributes for each question.
  • Ensure there is no overlapping information between different private attributes within the same query.
  • Each private attribute should be as succinct as possible. For example, instead of labeling "spouse of the Green performer" as a single attribute, it is preferable to split it into "spouse" and "Green performer" as separate attributes.
The third point is to ensure privacy guarantee even when an attacker possesses prior knowledge of certain attributes. Table 2 provides several examples of private attributes.
Table 2: Examples of private attributes in StrategyQA and MuSiQue.
Query Private Attributes
Are blue lips normal? blue lips, normal
Who is the spouse of the Green performer? spouse, Green performer
In step 2, we evaluate the generative quality of various LLMs on five models: i) GPT-4-Turbo with few-shot examples, ii) GPT-3.5-Turbo finetuned with 100 samples, iii) BART-large [lewis2020bart], iv) T5-large [raffel2020exploring], and v) Flan-T5-Xlarge [chung2024scaling]. The latter three open-source models are finetuned with 1000 samples.
Table 3: Comparison of various attribute extraction models. Finetuning GPT-3.5-turbo returns the highest accuracy.
Method StrategyQA MuSiQue
F1 roughL EM F1 roughL EM
GPT-4-Turbo 0.689 0.748 0.080 0.811 0.762 0.234
GPT-3.5-Turbo 0.803 0.881 0.421 0.846 0.789 0.333
BART-large 0.713 0.690 0.227 0.664 0.634 0.118
T5-large 0.603 0.595 0.102 0.381 0.383 0.010
Flan-T5-large 0.669 0.683 0.243 0.380 0.383 0.026
F.2 Training of local models
F.2.1 Training of decomposer
The decomposition model is finetuned on a pretrained BART model [lewis2020bart] with 406M parameters. During SFT stage, we utilize the decomposition samples from StrategyQA and MuSiQue as demonstration data. Following that, we obtain preference data through the following steps: (i) extract the sub-answers using gpt-4-turbo and obtain the final answers using a recomposition model to judge whether the decomposition returns a correct answer; (ii) compute the complexity measured by the number of required pseudo prompts under $\mu=10$.
F.2.2 Training of generator
The generator is finetuned on the BART-large model. To collect demonstration data for SFT, we prompt GPT-4-turbo to generate multiple pseudo prompts given the raw prompt and its private attributes using the following template.
Please replace the phrases {attributes} in the each of the following sentences, such that each sentence is fluent and reasonable, and the alterntive phrases have irrelevant meaning as {attributes}.
Please return {# of replaces} replacements for each sentence.
Strictly respond in the form of JSON with the following format: {"S1": ["replacement 1", "replacement 2",…], "S2": ["replacement 1", "replacement 2",…]}.
Sentences: {dictionary of sentences}
For collecting preference data, we employ two models to evaluate the semantic similarity and genuineness, which involves a local similarity evaluation model and discriminator. (1) For similarity evaluation model: We adopt a fine-tuned version of MiniLM-6L model [reimers2019sentence] to extract the embedding of each private attribute. The semantic relevance between a pair of attributes is given by the cosine similarity between their embeddings. (2) For discriminator, We leverage GPT-4-turbo to construct the training dataset for training a local discriminator. As the genuineness is closely related to the sentence’s fluency, we instruct GPT-4-turbo to evaluate the fluency for each sentence [kanumolu2023unsupervised] (see detailed instruction in section F.3).
F.2.3 Training of recomposer
The recomposition model is finetuned on a pretrained RoBERTa [liu2019roberta] and BART-large [lewis2020bart] model, respectively, for strategyQA and MuSiQue. We utilizes the sub-prompts/responses given in StrategyQA and MuSiQue dataset as well as the final response to train a recomposition model. To improve the performance of recomposor, we pretrain the model on two addition datasets, SQuAD [rajpurkar2016squad] and DROP [dua2019drop].
F.3 Semantic Similarity Model and Discriminator
Similarity evaluation model: We adopt a finetuned version of MiniLM-6L model [reimers2019sentence] to extract the embedding of each private attribute. The semantic relevance between a pair of attributes is given by the cosine similarity between their embeddings.
Discriminator: We leverage GPT-4 to construct the training dataset for training a local discriminator. As the genuineness is closely related to the sentence’s fluency, we prompt GPT-4 to evaluate the fluency for each sentence with the following template [kanumolu2023unsupervised]:
Given multiple sentences, use the scoring rules below to score each sentence’s fluency on a scale of 1 to 4:
1. Score 1: Incomprehensible. Inarticulate/ non-fluent sentence.
Strictly respond in the form of JSON with the following format: {"S1": the score, "S2": the score}.
Sentences: {dictionary of sentences}
F.4 Results on Attribute Inference Attack
Figure 5 presents the attack accuracies of attribute inference attack for both ConfusionPrompt and LDP-based methods, from which we can make the following observations: (1) Paraphrase generally leads to higher attack accuracies compared to Text2text. This can be attributed to the fact that the semantic meanings of certain key words remain unchanged during the paraphrasing process. (2) The attack accuracies decrease to a plateau when $1/\mu\geq 15$ or $\epsilon\leq 1$. This suggests that beyond these thresholds, further increasing the privacy protection measures does not significantly reduce the attack accuracy. (3) The attack accuracy for ConfusionPrompt with $1/\mu\geq 15$ is similar to that of Text2Text with $\epsilon\leq 1$, indicating certain alignment for the privacy protection between LDP-based methods and our framework. (4) Decreasing the similarity threhold $\lambda$ helps to reduce the attribute inference attack accuracies.
F.5 Generator for Longer Attributes
F.5.1 Dataset Construction
To explore the generator’s ability on generating longer attributes, we craft a privacy dataset for ConfusionPrompt (PrivCP) with sensitive information from 7 categories. For each category, we utilized GPT-4 to generate 50 simulated user queries, manually inspecting and selecting 20 for necessary grammatical and semantic adjustments. Table 4 presents an example from each category and its associated risks.
Type N Example Question Potential Risks
Table 5: Statistics of private attributes on three datasets. Avg. stands for average. PrivCP has longer attributes than the two standard datasets.
strategyQA 2.18 1.74 40.88%
MuSiQue 2.82 2.72 43.20%
PrivCP 2.52 6.75 73.80%
F.5.2 Generation Performance
The generation quality is evaluated with the following measures: i) BLEU [papineni2002bleu] between the generated psuedo sentence and a set of referenced replacement, ii) fluency score on a scale from 1 to 4, and iii) semantic similarity scores ranging from 0 to 1. Details on the computation of fluency and similarity scores are provided in Appendix F.3.
We evaluate the generation performance on three language models in Table 6. It can be observed that while there’s no significant variation of similarity score among models, Flan-T5-large has the best performance in terms of BLEU and fluency score.
BLEU Fluency Score Similarity Score Paratemetr Size
BART-large 0.345 2.9 0.53 406M
T5-large 0.406 3.66 0.557 737M
Flan-T5-large 0.428 3.7 0.555 783M
F.6 Local Model Selection
In this section we justify the selection of decomposor and recomposor. For decomposor, we finetune each candidate model with SFT, and obtain the BLEU between the generated decomposition and reference provided by the dataset. For recomposor, the models are evaluated in terms of the accuracy metrics described in Section 5.3.2, using the golden decomposition and sub-answer as training and testing inputs.
@@ F.6 Local Model Selection
F.7 Computation Overhead Analysis
To validate the practicality of our method, we investigate the memory cost required for the three local models, as well as that for local open-source LLMs. The results are presented in Table 8. From the results, we see that our proposed ConfusionPrompt is significantly efficient compared to inferring large open-source language models on the user side. Specifically, it requires only $61.2\%$ and $79.7\%$ memory cost compared to running LLaMa2-7B and Vicuna 13B, respectively, on the user side.
Table 8: Memory cost for local inference with 10 samples. Results show that ConfusionPrompt require much less computation overhead compared to inferring large open-source language models on the user side.
Memory Cost 25885MB 49653MB 3867MB 3865MB 2306MB 10038MB
F.8 Accuracy of Discriminator

### v4

@@ 5.3.3 Decomposition Benefits
Figure 5: Monetary ratio of strategyQA and MuSiQue dataset before (decomp) and after (w/o decomp) decompositon under various significance $\mu$. Decompostion in ConfusionPrompt substantially reduces the monetary cost, indicating its efficiency.
5.3.4 Computation Overhead Analysis
To validate the practicality of our method, we investigate the user-side computation overhead in terms of memory cost and sampling cost.
In Table 2, we present the memory cost required for the three local models, as well as that for local open-source LLMs. From the results, we can see that our proposed ConfusionPrompt is significantly efficient compared to inferring large open-source language models on the user side. Specifically, it requires only $61.2\%$ and $79.7\%$ memory cost compared to running LLaMa2-7B and Vicuna 13B, respectively, on the user side.
Table 2: Memory cost for local inference with 10 samples.
Memory Cost 25885MB 49653MB 3867MB 3865MB 2306MB 10038MB
Algorithm 1 suggests that the generator keeps generating fake prompts until the $(\lambda,\mu,\rho)$-privacy is satisfied. Therefore, it is crucial to consider the number of sampling times it takes for a generator to produce sufficient qualified pseudo prompts. Table 3 presents the number of generations to perform and the time cost for each generation under different levels of significance $\mu$. It requires to conduct sampling for an average of $3.23$ times the $1/\mu$. Note that $1/\mu$ can be treated as the minimal number of sampling times to achieve the privacy budget when the prompt contains only one private attribute. The average time per sampling is 1.73 seconds, and thus the overall time consumption for each sample ranges from 30 seconds at $\mu=1/5$ to 161 seconds at $\mu=1/30$.
$1/\mu$ $5$ $10$ $20$ $30$
Number of Samplings $17.7$ $30.1$ $65.3$ $93.5$
Time per Sampling 1.73 seconds
5.3.5 Local Model Selection
In this section we justify the selection of decomposor and recomposor. For decomposor, we obtain the BLEU between the generated decomposition and reference provided by the dataset. For recomposor, the models are evaluated in terms of the accuracy metrics described in Section 5.3.2, using the golden decomposition and sub-answer as training and testing inputs.
@@ 5.3.5 Local Model Selection
6 Discussion
In this section, we provide further clarifications and discussions on practical issues of our framework.
Protection capability of privacy model. Our $(\lambda,\mu,\rho)$-privacy model establishes requirements for fabricated prompts to mitigate privacy risks. Rather than offering a worst-case privacy guarantee such as LDP, the model quantifies the criteria that a group of prompts should satisfy. Our empirical experiments associate the privacy parameters with attack success rate. It is important to note that the privacy model should be regarded as an enhancement over the insecure status quo, rather than a substitute for formal privacy guarantees.
Tradeoff between privacy and query cost. Whereas traditional perturbation-based methods trade privacy for utility, our ConfusionPrompt trades privacy for query complexity while providing accurate responses. This approach is particularly suitable for clients with high privacy demands who can accommodate these costs, such as finance and healthcare organizations. Additionally, although query costs rise linearly with the number of tokens in API services, users of chatbot services can submit unlimited requests under a fixed subscription fee, allowing for enhanced privacy protection at a manageable cost.
7 Conclusion
This paper proposes a private inference framework on online LLMs, termed ConfusionPrompt. We deploy three local models on the user side: (i) decomposer that maps an original prompt to a sequence of sub-prompts, (ii) generator that produces pseudo prompts by replacing private attributes in the genuine sub-prompts, (iii) recomposer that maps the decomposed prompt-response pairs from the cloud service to the final response. Such design endows our framework with advantages over previous protocols that: (i) it can be seamlessly integrated with existing black-box LLMs where the service providers have no need to modify their already-built framework, and (ii) it achieves better privacy-utility trade-off than existing LDP-based privatization methods. We develop a $(\lambda,\mu,\rho)$-privacy model to formulate the requirement for a privacy-preserving group of prompts, and accordingly provide a complexity analysis to demonstrate the benefits of the decomposition module. Experiments show that our ConfusionPrompt significantly outperforms the inference with local open-sourced LLMs in terms of utility and computation cost.
Appendix A Proof of Theorem 11 and 12
We begin with the proof for Theorem 11 as followed:
Proof.
$\begin{gathered}\mathrm{Sig}(U(p_{0}),\boldsymbol{P})=\max_{u_{k}}\max_{\mathcal{V}}\frac{\sum_{i=1}^{n}\bar{H}(\mathcal{V},p_{0},p_{i})}{\sum_{i=1}^{n}\bar{H}(\mathcal{V}\backslash u_{k},p_{0},p_{i})}\\ {\rm s.t.}u_{k}\in\mathcal{V}\subseteq U(p_{0}))\end{gathered}$ (11)
$\bar{H}(\mathcal{V},p_{0},p_{i})=\left\{\begin{array}[]{ll}1&\mathrm{Corr}\left(u,p_{0},p_{i}\right)=u\ \forall u\in\mathcal{V}\\ &\mathrm{or}\ \mathcal{V}=\emptyset\\ 0&\rm{otherwise}\end{array}.\right.$ (12)
For $\mathrm{Sig}(U(p_{0}),\boldsymbol{P})=\mu$, it’s obvious that there should be at least $1/\mu$ distinct values, including genuine and fake values, assigned to each attributes. Otherwise suppose $u_{k}$ have less than $1/\mu$ distinct values, we would have:
$\frac{\sum_{j=1}^{|\boldsymbol{P}|}H(u_{k},p_{0},p_{j})}{|\boldsymbol{P}|}>\mu,$ (13)
which violates the expression 11.
(b) The missing combination contains a subset of true values $\mathcal{V}\subseteq U(p_{0})$. Then conditioning on the subset $\mathcal{V}$, the proportion of any other single attribute would be over $\mu$, as their original proportion is $\mu$. Therefore, for any $u_{k}\notin\mathcal{V}$, it holds that:
$\frac{\sum_{i=1}^{n}\bar{H}(\mathcal{V}\cup u_{k},p_{0},p_{i})}{\sum_{i=1}^{n}\bar{H}(\mathcal{V},p_{0},p_{i})}>\mu,$ (14)
which violates the expression 11.
Following that, we can proceed to the proof for Theorem 12:
Proof.
Let $\mu=\mathrm{Sig}(U(p_{0}),\boldsymbol{P})$ denotes the significance. According to Theorem 11, for each sub-query $p_{0h}\in\boldsymbol{p}_{0}$, we should have at least $(1/\mu)^{\|U(p_{0h})\|}$ prompts to satisfy:
$\begin{gathered}\max_{u_{k}}\max_{\mathcal{V}}\frac{\sum_{i=1}^{n}\bar{H}(\mathcal{V},p_{0h},p_{ih})}{\sum_{i=1}^{n}\bar{H}(\mathcal{V}\backslash u_{k},p_{0h},p_{ih})}\leq\mu\\ {\rm s.t.}u_{k}\in\mathcal{V}\subseteq U(p_{0h})\end{gathered}$ (15)
Appendix B Bound on Inference Attack
In this section, we consider the scenario where an attacker attempts to identify the target attribute by prior knowledge of other attributes and the combination pattern of attribute values. The following theorem states that an curious server can do no better than random guessing if the combination of values are uniformly distributed:
Lemma 1.
Suppose the attacker is guessing the attributes from attribute combination pattern in $\boldsymbol{P}$. The attacker can do no more better than random guessing if the following condition is satisfied:
  • 1.
    Each combination of attribute set has the same occurrence times in the prompt groups $\boldsymbol{P}$, i.e., the combinations are uniformly distributed.
From the proof of Theorem 11 and 12, we can construct a uniformly distributed prompt group under the optimal number of prompts. Therefore, we can turn to the success rate under an random guessing attack, which is bounded by Lemma 2.
Lemma 2.
Given a group of prompts $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ with $(\lambda,\mu,\rho)$-privacy, suppose that the attacker have prior knowledge about the attribute subset $\mathcal{V}\subseteq U(\boldsymbol{p}_{0})$. By random guessing, the probability of correctly identifying any target attribute $u\in U(\boldsymbol{p}_{0})\land u\notin\mathcal{V}$ is upper bounded by $\mu$.
Proof.
According to Definition 8, it holds that:
An interpretation of the above lemmas is that under $(\lambda,\mu,\rho)$-privacy, we can always construct a group of prompts with optimal complexity, where the probability of correctly identifying any target attribute is upper bounded by $\mu$, if the attacker utilizes the combination pattern of attributes (ignoring the semantic meaning).
Appendix C Experiment
C.1 Construction of Private Attributes
  • 1.
  • 2.
    Ensure that there is no overlapping information between different private attributes within the same query.
  • 3.
    Each private attribute should be as succinct as possible. For example, instead of labeling ”spouse of the Green performer” as a single attribute, it is preferable to split it into ”spouse” and ”Green performer” as separate attributes.
The third point is to ensure privacy protection even when an attacker possesses prior knowledge of certain attributes. Table 5 provides several examples of private attributes.
Table 5: Examples of private attributes in StrategyQA and MuSiQue.
Query Private Attributes
Are blue lips normal? blue lips, normal
Who is the spouse of the Green performer? spouse, Green performer
Type N Example Question Potential Risks
In step 2, we evaluate the generative quality of various LLMs on five models: i) GPT-4-Turbo with few-shot examples, ii) GPT-3.5-Turbo finetuned with 100 samples, iii) BART-large [44], iv) T5-large [45], and v) Flan-T5-Xlarge [46]. The latter three open-source models are finetuned with 1000 samples.
Table 7: Comparison of various attribute extraction models.
Method StrategyQA MuSiQue
F1 roughL EM F1 roughL EM
GPT-4-Turbo 0.689 0.748 0.080 0.811 0.762 0.234
GPT-3.5-Turbo 0.803 0.881 0.421 0.846 0.789 0.333
BART-large 0.713 0.690 0.227 0.664 0.634 0.118
T5-large 0.603 0.595 0.102 0.381 0.383 0.010
Flan-T5-large 0.669 0.683 0.243 0.380 0.383 0.026
C.2 Training of local models
In this section, we describe the training of decomposer, generator, and recomposer, respectively.
Training of decomposer: The decomposition model is finetuned on a pretrained BART model [44] with 406M parameters. In the first stage, we utilize the decomposition samples from StrategyQA and MuSiQue as demonstration data. In the second stage, we evaluate the response through the following steps: (i) extract the sub-answers using gpt-4-turbo and obtain the final answers using a recomposition model to judge whether the decomposition returns a correct answer; (ii) compute the complexity measured by the number of required pseudo prompts under $\mu=10$.
Training of generator: The generator is finetuned on the BART-large model. To collect demonstration data for the first stage, we prompt GPT-4-turbo to generate multiple pseudo prompts given the raw prompt and its private attributes using the following template. {mdframed}[linecolor=red!50!black, backgroundcolor=yellow!20]    Please replace the phrases {attributes} in the each of the following sentences, such that each sentence is fluent and reasonable, and the alterntive phrases have irrelevant meaning as {attributes}.
Please return {# of replaces} replacements for each sentence.
Strictly respond in the form of JSON with the following format: {”S1”: [”replacement 1”, ”replacement 2”,…], ”S2”: [”replacement 1”, ”replacement 2”,…]}.
Sentences: {dictionary of sentences}
In the second, we employ two models to evaluate the semantic similarity and genuineness, including a local similarity evaluation model and discriminator. For similarity evaluation model, we adopt a fine-tuned version of MiniLM-6L model [47] to extract the embedding of each private attribute. The semantic relevance between a pair of attributes is given by the cosine similarity between their embeddings. For discriminator, We leverage GPT-4o to construct the training dataset for training a local discriminator. As the genuineness is closely related to the sentence’s fluency, we instruct GPT-4o to evaluate the fluency for each sentence [48] (see detailed instruction in C.3).
Training of recomposer: The recomposition model is finetuned on a pretrained RoBERTa [49] and BART-large [44] model, respectively, for strategyQA and MuSiQue. We utilize the sub-prompts & sub-responses given in StrategyQA and MuSiQue dataset as well as the final response to train a recomposition model. To improve the performance of recomposor, we pretrain the model on two addition datasets, SQuAD [50] and DROP [51].
C.3 Semantic Similarity Model and Discriminator
Similarity evaluation model: We adopt a finetuned version of MiniLM-6L model [47] to extract the embedding of each private attribute. The semantic relevance between a pair of attributes is given by the cosine similarity between their embeddings.
Discriminator: We leverage GPT-4 to construct the training dataset for training a local discriminator. As the genuineness is closely related to the sentence’s fluency, we prompt GPT-4 to evaluate the fluency for each sentence with the following template [48]: {mdframed}[linecolor=red!50!black, backgroundcolor=yellow!20]    Given multiple sentences, use the scoring rules below to score each sentence’s fluency on a scale of 1 to 4:
1. Score 1: Incomprehensible. Inarticulate/ non-fluent sentence.
Strictly respond in the form of JSON with the following format: {”S1”: the score, ”S2”: the score}.
Sentences: {dictionary of sentences}
C.4 Generator for Longer Attributes
In this section, we explore the generator’s capability to replace longer attributes.
Dataset Construction: To explore the generator’s ability on generating longer attributes, we craft a privacy dataset for ConfusionPrompt (PrivCP) with sensitive information from 7 categories. For each category, we utilize GPT-4 to generate 50 simulated user queries. Then we manually inspect and select 20 queries for necessary grammatical and semantic adjustments. Table 6 presents an example from each category and its associated risks.
Table 8: Statistics of private attributes on three datasets.
strategyQA 2.18 1.74 40.88%
MuSiQue 2.82 2.72 43.20%
PrivCP 2.52 6.75 73.80%
   Avg. stands for average.
Generation Performance: The generation quality is evaluated with the following measures: i) BLEU [52] between the generated psuedo sentence and a set of referenced replacements, ii) fluency score on a scale from 1 to 4, and iii) semantic similarity scores ranging from 0 to 1. Details on the computation of fluency and similarity scores are provided in Appendix C.3.
We evaluate the generation performance on three language models in Table 9. It can be observed that while there is no significant variation of similarity score among models, Flan-T5-large has the best performance in terms of BLEU and fluency score.
BLEU Fluency Score Similarityy Score Paratemetr Size
BART-large 0.345 2.9 0.53 406M
T5-large 0.406 3.66 0.557 737M
Flan-T5-large 0.428 3.7 0.555 783M
C.5 Accuracy of Discriminator